westcoast wholesalers expects to sell approximately 3000 boxes of soap during 2018. the cost of placing an order for soap is estimated to be R18. the inventory holding of one box of soap is R2. Calculate the annual economic order quantity for the soap


1
Expert's answer
2021-06-09T08:59:33-0400

The annual economic order quantity for the soap is:

"EOQ = \sqrt{\frac{2*3,000*18}{2}} = 232.38."